OpenStudy (anonymous):

What is the quotient: (72y^3 + 12y^2 – 30y) ÷ 6y ?

OpenStudy (anonymous):

12y²+2y-5 ?

OpenStudy (mathmate):

When a polynomial is divided by a monomial, you only have to divide each term by the monomial. For example, 72y^3/6y=12y^2, +12y^2/6y=2y -30y/6y=-5 Can you put the answer together?

OpenStudy (anonymous):

(6y-5)(2y+1)